If 41.6g of chromium is fully reacted with oxygen to produce 60.8g of chromium oxide, what is the empirical formula of the chromium oxide
Answer #1

Mass of Cr2O3 =151.99 g/mol

We know atomic mass of oxygen =16g

Mass of oxygen in chromium oxide =(16×3/151.99 )×60.8

=19.20g

Percentage of Chromium= (41.6/41.6+19.20)×100%

=68.42%

Percentage of oxygen = (19.2/41.6+19.2)× 100%

=31.57%

Element Percentage of element atomic mass moles of atom mole ratio
Chromium 68.42 52 1.315 1
Oxygen 31.57 16 1.973 1.5

Now rounding off to simplest whole no ratio

Chromium =2

Oxygen= 3

Therefore empirical formula = Cr2​​​​​​O3